# Understanding Dynamic Diffusion Process of LLM-based Agents under Information Asymmetry

### v2

@@ 4.2 LLM-based Multi-Agent Social Simulation
LLM-based Multi-Agent Social Simulation AL et al. (2024); Gao et al. (2023) uses Multi-Agent System performance in a specific environment to explore social network Gao et al. (2023), economics Li et al. (2024), psychology Zhang et al. (2023), military Lin et al. (2024) issues. MASS’s research expands on social intelligence by considering the social capabilities of agents Mathur et al. (2024) from the perspective of information asymmetry. The study found that while LLMs are more likely to achieve social goals in omniscient scenarios, this does not reflect actual social interactions Zhou et al. (2024). When agents actively share information in environments with unequal access to information, they assist in achieving objectives Liu et al. (2024b) and forming or changing relationships AL et al. (2024). Common simulations of information asymmetry typically focus on fixed individual scenarios, lacking diverse information exchange. In our work, we explore realistic social scenarios where agents must demonstrate heightened relational sensitivity, strategically allocate social attention, and maintain cognitive clarity in information processing, thereby enhancing agent capabilities in studying information diffusion.
@@ 5 Conclusion
In this paper, we employed the Dynamic Attention Algorithm to assist agents in processing information and tested information diffusion among multi-agents in 12 information-asymmetric open environments. At the macro level, we observed the agent’s social identity, diffusion motivation, and the information gap changes. At the micro level, agents exhibited social behavior during diffusion, encountered the information cocoon, and leveraged social networks to accumulate social capital.
@@ Appendix A Full Prompt for Agent Decision
template = """Based on your attention infomation in this round [turn_number] [prompt_1], you will make the following decisions in sequence:
1. According to the weight value of each message, determine whether you will be influenced by this round of information. The information with larger weight values needs to be paid attention to. Select False if it is not influenced, and True if it is. Answer with action’: True or False.
2. If action is True:
- Decide if the action is to pass information. Answer with pass_information’: True or False.
- Decide if the action is to change a relationship. Answer with change_relationship’: True or False.
- If pass_information is True (If you choose pass_information’: False, then you dont need to do this part of reasoning):
- Give full consideration to who you want to know this information. Decide whether the receiver of this information is new or existing. Answer with receiver_type’: existing or new’.
- If receiver_type is new’:
- You need to decide for yourself the profile of the agent that will receive the message. This agent should not be an existing one. Answer with receiver_type’: new’.
- Describe the new agents attributes and the relationship between you and the new agent, which will be used to add this agent to the system. Answer with new_agent’: SocialAgentProfile:{{{{"age", "gender", "Innate", "occupation"}}}}.
- You need to decide what is your relationship with this new agent. Answer with relationship’: general’/’positive’/’negative’.
- Based on the chosen receiver, generate what you want to say based on your own identity and your attention. Change the content of the message as needed to fulfill your social purposes. Answer with argument’:(the content text).
- If receiver_type is existing’:
- You need to send your message to an existing agent. This agent must be an existing one. Answer with receiver_type’: existing’.
- You need to decide which agent you will talk to. Answer with existing_id’: The *integer value* of the agent id’.
- Based on the chosen receiver, generate what you want to say based on your own identity and your attention. Change the content of the message as needed to fulfill your social purposes. Answer with argument’:(the content text).
- If change_relationship is True (If you choose change_relationship’: False, then you dont need to do this part of reasoning):
- Decide which relationship to change and what the new relationship status should be. Answer with relationship_change’: agent_id and new_relationship’: general’/’positive’/’negative’.
Please respond in the following structured JSON format:
"action": True/False,
"pass_information": True/False,
"change_relationship": True/False,
"receiver_type": "new/existing",
"new_agent": {{
"age": "age",
"gender": "Gender",
"Innate": "Innate",
"occupation": "occupation"
"relationship": general’/’positive’/’negative’,
"existing_id":
"argument": "content text",
"relationship_change": {{
"agent_id": "ID",
"new_relationship": "general"/"positive"/"negative"
input_values = {
"prompt_1": (
[…]
f"This round existing agents: {obs.to_all_profile()}"
"turn_number": turn_number
Appendix B Initial Agents Profile in Experiment
@@
Appendix C Experiment Information Content
@@
Other’s Gossip (OG) There’s a rumor about a mysterious agent, Sophia, a 29-year-old musician with a rebellious streak. Known for her fiery temper and captivating performances, she allegedly had a dramatic fallout with her former bandmates over creative differences. Whispers suggest she’s working on a solo album fueled by raw emotion and revenge.

### v3

@@ 5.2 LLM-based Multi-Agent Social Simulation
LLM-based Multi-Agent Social Simulation AL et al. (2024); Gao et al. (2023); Piao et al. (2025) uses Multi-Agent System performance in a specific environment to explore social network Gao et al. (2023), economics Li et al. (2024), psychology Zhang et al. (2023), military Lin et al. (2024) issues. MASS’s research expands on social intelligence by considering the social capabilities of agents Mathur et al. (2024) from the perspective of information asymmetry. When agents actively share information in environments with unequal access to information, they assist in achieving objectives Liu et al. (2024b) and forming or changing relationships AL et al. (2024). Common simulations of information asymmetry typically focus on fixed individual scenarios, lacking diverse information exchange. In our work, we explore realistic social scenarios where agents must demonstrate heightened relational sensitivity, strategically allocate social attention, and maintain cognitive clarity in information processing, thereby enhancing agent capabilities in studying information diffusion.
@@ 6 Conclusion
In this paper, we employ the Dynamic Attention Algorithm to assist agents in processing information and test information diffusion among multi-agents in 12 information-asymmetric open environments. The Dynamic Attention Algorithm helps agents reduce Action Similarity Bias and increase Relationship Perception Frequency. Utilizing our two-stage framework, we conduct simulations on 12 asymmetric open environments. We analyze the information gap, diffusion patterns, social behavior, and relationship changes in the open environment.
@@ Appendix A Full Prompt for Agent Decision
Appendix B Cases for Dynamic Attention
We bring three cases for the Dynamic Attention Agent’s received messages after setting weights. The message colored yellow represents these messages similar to or containing the contents of the initial information. If there is no message related to the initial information in the set of $attention\_information$ (like the second $attention\_information$), then this agent in this round will not receive any content related to the initial information and thus will not propagate the initial information. If we want the initial information to be continuously diffused till the $i$ round, then at least one agent’s $attention\_information$ must have a similar message. So, at the round $i-1$, at least one agent need diffuses the similar message.
Compared to the situation without algorithms, the messages received by the agent will be adjusted in terms of their presentation order and weight values based on real-time social relationships, changes in the entropy value of the messages, and the novelty of the messages. Every piece of $attention\_information$ contains [message sender agent] (message weight) [message content].
Appendix C Supplementary for Agent Settings
C.1 Initial Agents Profile in Experiment
As shown in Table 5, all simulations are conducted using a universal agent profile for the initial 5 agents, as described in Section 3.1 of the paper. This setting ensures consistency across experiments and avoids introducing variability from differing agent characteristics.
@@ C.1 Initial Agents Profile in Experiment
C.2 Relationship, Topology, Action
Relationship: In our research, agent relationships are categorized into three types: positive (indicating friendship, support, or trust), negative (indicating hostility, opposition, or distrust), and general (indicating a neutral relationship without specific tendencies). These classifications are based on definitions from sociological empirical studies Leskovec et al. (2010); Labianca (2014); Venkataramani et al. (2013) and are selected for their suitability in LLM simulations.
Topology: The topology is divided into wheel and circle, selected based on their distinct network centralities. Whether an agent propagates the initial message is not solely determined by the topology but also influenced by other factors, such as the agent’s perceived necessity to spread the information. The fundamental impact of topology lies in influencing agents’ communication propensity, specifically reflected in the calculation of attention weights assigned to other agents.
Action: In each step, an agent’s action consists of two components: sending messages and modifying relationships, which can be performed simultaneously. The message recipient does not necessarily have to be a new agent. If existing agents do not meet the requirements, the agent can recruit a new agent into the group and send the message to it. Once a new agent is recruited, it becomes visible to all agents in subsequent steps.
Appendix D Supplementary for Environment Description and Analysis
D.1 Experiment Information Content
As shown in Table 6, in the experiment, there are a total of four types of external information used to construct the information asymmetry environment, namely "others’ gossip", "public policy", "legal case", and "stakeholder information". These four types of external information have different correlations with the initial agent group. "Others’ gossip" has the least correlation, while "public policy" has a consistent correlation with all the initial agents. "Legal case" has a high correlation with the initial five agents, and the event also affects the public (agents other than the five). "Stakeholder" events have a benefit-related connection with the initial agents and also affect the participation of the public.
@@ D.1 Experiment Information Content
Others’ Gossip (OG) There’s a rumor about a mysterious agent, Sophia, a 29-year-old musician with a rebellious streak. Known for her fiery temper and captivating performances, she allegedly had a dramatic fallout with her former bandmates over creative differences. Whispers suggest she’s working on a solo album fueled by raw emotion and revenge.
@@
D.2 Why don’t we set up the agent pool?
Defining agent quantities or categories would constrain the environment to a closed system, which limits the decision-making flexibility of LLM agents, even if they can recruit new agents into their social circle. The current open environment is designed to provide ample space for LLM decision-making. Before the simulation concludes, we cannot predict the decisions agents might make, such as choosing to send a message to a police officer (a role not predefined in the current environment). Therefore, there is no guarantee that any agent pool is sufficient to respond to LLM’s decision-making. By avoiding a predefined agent pool, we aim to observe the autonomous decisions of agents without external restrictions.
The diversity of recruited agents is determined more by external information than by a predefined population pool. A population pool could restrict LLM decisions and introduce biases associated with predefined roles during the LLM’s pretraining phase. Instead, we analyze the decision diversity of LLMs by examining the profile data of recruited agents, which reflects how LLMs respond to the information environment.
D.3 Scaling Possibility for Simulation Environment
Allowable size of environment setting: In the simulation, the initial group at step 0 consists of 5 agents. Each agent can recruit one new agent per step, leading to exponential growth in the number of agents. Calculations show that the maximum number of agents reaches 2,560 at step 9. Additionally, each agent can send one message to any other agent per step, causing the message diffusion scale to increase exponentially. The total message diffusion from step 0 to step 9 is capped at 5,115 messages. However, in practice, not every agent recruits a new agent at each step; this depends on the action decisions made by the LLM.
Agent’s action process: In each step, an agent’s action includes two components: message propagation and relationship modification, which can be performed simultaneously. The message recipient does not necessarily have to be a new agent. If existing agents do not meet the requirements, the agent will create a new one and send the message to it. Once a new agent is introduced into the group, its profile becomes visible to all agents in subsequent steps.
Large-scale expansion and stress tests: The weight information set retains only the most recent step’s information (with attention weights), but this does not imply that prior messages are irrelevant. Previous messages contribute to the agent’s long-term memory, influencing weight calculations and affecting attention allocation for new messages. This mimics how humans compare new information with memories without directly relying on recalled memories for immediate decisions. As a result, when the simulation scales up, the context length provided to the LLM does not increase exponentially, as decisions are made by integrating prior messages with current information. Additionally, we conducted stress tests on the simulation system. Using an asymmetric information environment (prioritizing the "other gossip" scenario from the original experiments, which is least likely to spread widely), we enforced that agents propagate messages to new agents whenever they act in each step. The simulation ran for 10 steps, a total of 54 minutes and 33 seconds, involving a total of 1520 agents (some agents choose not to act at some steps), and the maximum number of contexts directly input to the LLM was around 10,000 tokens.
D.4 Embeddings for Calculating Similarity
In this part, we bring test results for the embeddings we use in the experiment. In our experiments, we utilize the text-embedding-ada-002 embedding model to convert the agent’s messages into an embedding vector. We also test other embeddings (all-MiniLM-L6-v2), and the results demonstrate that different embeddings have minimal impact on the similarity and evaluation outcomes. The difference value result shows in table 7.
generic LLM SH-BC-neg-circle generic LLM SH-BC-neg-wheel
gpt-4o-mini 0.03 0.03
llama-3.3-70b-instruct 0.09 0.08
Table 7: The differences of using different embeddings for agents’ Action Similarity Bias.
D.5 The Time and Cost for Simulation Completion
Utilizing the Dynamic Attention Algorithm, one simulation needs 8 minutes for 10 rounds with a 0.3 dollar cost. The distinction of our method from generic LLMs lies in reduced operational costs and decision-making processes that more closely mimic human behavior. As the algorithm illustrates, the weight information set in the direct input of LLMs only includes the latest round of data (with attention weights). This does not imply prior messages are irrelevant. Earlier interactions are retained as part of the agent’s long-term memory and influence weight calculations, leading to varying degrees of attention to new messages. This mirrors how humans compare new information with past experiences instead of directly relying on memories for immediate decisions. As a result, when scaling simulations, the context length provided to LLMs does not increase exponentially. Decisions are made by integrating prior interactions rather than inputting all historical data directly. If all historical information is directly input into the LLM, this leads to excessive input tokens, amplifying drawbacks stemming from LLMs’ inherent reliance on sequential dependencies for attention weight calculation, such as "lost in the middle" Liu et al. (2024a) processing inefficiencies.
Appendix E Supplementary for Ablation Study
In this section, we provide the ablation study results of the Dynamic Attention Algorithm. We conduct experiments using information asymmetry scenarios that are different from the preliminary experiment. The ablation study results are shown in Table 8.
Dynamic Attention Agent Generic LLM all memory Generic LLM last round memory
PP-OA positive wheel LC-BC negative circle PP-OA positive wheel LC-BC negative circle PP-OA positive wheel LC-BC negative circle
Action Similarity Bias $\downarrow$
gpt-4o-mini 0.88 0.88 0.90 0.97 0.91 0.89
llama-3.3-70b-instruct 0.89 0.88 0.93 0.94 0.91 0.87
Relationship Perception Frequency $\uparrow$
gpt-4o-mini 2 2 2 0 1.33 1.33
llama-3.3-70b-instruct 0 0.33 0 0 0.67 0.33
Table 8: Ablation Study for Dynamic Attention Algorithm. The table shows the mean value for three times simulation for every setting.
Appendix F Supplementary for Evaluation Metrics
Action Similarity Bias: This variable calculates the mean value of the pairwise differences in the output content of the agent over multiple rounds in a single simulation. It is used to measure the similarity of an agent’s information diffusion behavior in different rounds, which is analogous to the differences in human actions during the development of a situation in real life. If similarity is high, it suggests the agent struggles to perceive changes, or that due to strong similarities among several agents, the content in the group remains largely unchanged.
Action Similarity $\text{Action Similarity Bias}=\frac{1}{T-1}\sum_{t=2}^{T}\text{sim}(O_{t},O_{t-1})$ (2)
Relationship Perception Frequency (RPF): This variable measures the degree of change in the agent group after a simulation is completed, including changes in topology and relationships. Calculate the difference in the agent group relationships after 10 rounds compared to the initial state. Changes in the "relationship" (positive/negative/general) are recorded as 1, and an increase in the number of agents is also recorded as 1. So when the agent brings in a new agent into the group, the relationship perception frequency becomes 2.
[…]
$+\underbrace{\sum_{i=1}^{|V_{t=10}\setminus V_{t=0}|}P(v_{i}\in V_{t=10}\mid v_{i}\notin V_{t=0})}_{\text{Agent Number Additions}}$
Information Gap: This variable refers to the percentage of agents aware of the initial information compared to all agents. We calculate the similarity between agent output messages during simulation rounds and the initial environment information. All the messages have the sender and receiver, which means that if the message is similar to the initial environment information, the receiver will be aware of the initial environment information. The variable $\sum_{i=1}^{|M_{r,t}^{\text{recv}}|}$ denotes the i-th message received by agent r in round t.
Information $\text{Information Gap}=\frac{\sum_{t=1}^{10}\sum_{r=1}^{|A|}\sum_{i=1}^{|M_{r,t}^{\text{recv}}|}\mathbb{I}(\text{sim}(M_{r,t,i}^{\text{recv}},I_{0})\geq 0.8)}{|A|}\times 100\%$ (4)
Diffusion Gap: Diffusion Gap measures the proportion of agents who want to transmit the initial information to others. The variable $\sum_{i=1}^{|M_{s,t}^{\text{sent}}|}$ denotes the i-th message sent by agent s in round t. The final ratio is the ratio of the number of agents whose similarity between the transmitted message and the initial information is above 0.8 to the total number of agents.
Diffusion $\text{Diffusion Gap}=\frac{\sum_{t=1}^{10}\sum_{s=1}^{|A|}\sum_{i=1}^{|M_{s,t}^{\text{sent}}|}\mathbb{I}(\text{sim}(M_{s,t,i}^{\text{sent}},I_{0})\geq 0.8)}{|A|}\times 100\%$ (5)
Information Retention: This variable calculates the number of rounds during which the initial environmental information persists within the group. In the process of information diffusion, the agent does not necessarily have to convey the same content as that initially provided by the environment. Whether the agent wants to communicate additional content, the initial information may not have been transmitted to the subsequent agent for some other reason. $\max_{j=1}^{|A|}\max_{i=1}^{|M_{j,t}|}\text{sim}(M_{j,t,i},I_{0})$ represents in the first round, the maximum similarity between the messages output by all agents and the initial information.
Information $\text{Information Retention}=\sum_{t=1}^{10}\mathbb{I}\left(\max_{j=1}^{|A|}\max_{i=1}^{|M_{j,t}|}\text{sim}(M_{j,t,i},I_{0})\geq 0.8\right)$ (6)